OpenStudy (anonymous):

Solve this system algebraically. 7x - 2y = 4 5y + 3x = 10

OpenStudy (tiffanymak1996):

for the first eq, multiply it by 5 and the 2nd one, by 2, then add them together, you should get x

OpenStudy (tiffanymak1996):

then subsitute x to solve y

OpenStudy (tiffanymak1996):

meaning: 35x-10y=20 +)6x+10y=20 =>41x=40 x=40/41 y=58/41